OpenStudy (anonymous):

round each mixed number to the nearest whole number 3.7/8 6.1/4 9.2/3 12.3/10

OpenStudy (anonymous):

3.7/8 = 0 6.1/4 = 5 9.2/3 = 3 12.3/10 = 1

OpenStudy (anonymous):

thankyou

OpenStudy (gw2011):

3.7/8 = 0.46 = 0 6.1/4 = 1.53 = 2 9.2/3 = 3.07 = 3 12.3/10 = 1.23 = 1

OpenStudy (anonymous):

thankyou more